OpenStudy (anonymous):

x^2+10x+61=0 solve the equation and complex solutions

OpenStudy (anonymous):

\[x^2+10x=-61\] \[(x+5)^2=-61+25=-36\] \[x+5=\pm\sqrt{-36}=\pm6i\] \[x=-5\pm6i\]

OpenStudy (anonymous):

the above method is rite